### 20 Dec 2023 v1

# Electron impact partial ionization cross sections: R-Carvone, 2-Butanol, Imidazole, and 2-Nitroimidazole

Abstract
In this work, we theoretically calculate the electron impact partial and total ionization cross sections of R-carvone , 2-butanol , imidazole and 2-nitroimidazole . We have used the Binary Encounter Bethe (BEB) model to obtain the total electron impact ionization cross sections (TICS). The modified BEB method in combination with mass spectrum data of the molecules is used to calculate the PICS of the cationic fragments originating from the parent molecule. Our PICS data for R-carvone and 2-butanol are in good agreement with the experimental data for all the cation fragments along with the TICS data. For imidazole and 2-nitroimidazole, the estimates of the PICS are reported for the first time in the present study. We have found that both the modified BEB method and the mass spectrum dependence method work effectively to estimate PICS if we have the information about the appearance energies and mass spectrum data of the target under investigation.
I Introduction:
Electron impact total ionization cross section (TICS) and partial ionization cross section (PICS) are important to many branches of sciences ranging from plasma physics to astrophysics and biological sciences Itikawa (2017); Kim et al. (2015); Boudaïffa et al. (2000). In medical sciences, radiation damage to DNA/RNA is a major concern, where it has been found that secondary electrons play a crucial role in single and double-strand breaks of DNA/RNA Boudaïffa et al. (2000); Sanche (2002). The electron impact cross sections data at low and intermediate energy are important inputs in Monte Carlo track (MCT) analysis, which is used to study damage to living cells using ionizing radiation Boudaïffa et al. (2000). In recent times, electron collision with biomolecules has geared up and many studies have been conducted both experimentally Rehman and Krishnakumar (2022a, b) and theoretically Gupta and Antony (2014); Gupta et al. (2014, 2019); Vinodkumar et al. (2006); Scheer et al. (2007) to provide more data for understanding the mechanism of damages caused due to the electrons. While performing the MCT simulations, only the single channel ionization is taken into account and the channels due to the dissociative ionization processes are not considered in the model, which may reduce the accuracy of the MCT analysis. Hence, the inclusion of dissociative ionization processes is important and will help us to make the MCT model more accurate. Hence, providing information on the PICS for the cations that are different due to their mass-to-charge ratio originating from their parent molecule is inevitable for accurate analysis Rehman and Krishnakumar (2022b). Even with the rapid surge of electron-molecule interaction studies, there are many important molecules for which data are scarce in the literature both experimental and theoretical. In the present study, we have investigated the PICS and TICS of R-carvone, 2-butanol, imidazole, and 2-nitroimidazole. Our motivation has been to provide the theoretical estimates of the PICS for all the cation fragments originating from the parent molecules using the recently developed methodsGraves et al. (2022); Goswami et al. (2022). Here we will highlight the importance and applications of each of these targets.
R-carvone is an organic molecule that belongs to the class of terpenoids and has a wide variety of applications in the food industry, agriculture, and pest control. R-carvone has a minty flavored odor, used in essential oils. A recent study has found this compound can be used in anti-fungal wrapper Boonruang et al. (2017); Soleimani et al. (2022) for packaging fruits which increases shelf life. Jones et al.performed an experimental and theoretical study on electron (e,2e) impact ionization dynamics on R-carvone and its enantiomer S-carvone Jones et al. (2019), using three-body distorted wave (3BDW) formalism for calculating the cross sections theoretically. There is a three-part study performed on this molecule by Lopes et al.Lopes et al. (2020a); Amorim et al. (2021a, b). In the first part, Lopes et al.Lopes et al. (2020a) presented the appearance energies for 35 cations and the electron impact mass spectrum in which they observed 103 peaks with the incident electron energy of 70 eV at the mass range of 1 amu till 151 amu and compared their findings with the data available at National Institute of Standards and Technology (NIST) web book Linstrom and Mallard (2001) and Spectral database for organic compounds (SDBS) SDB . Also, they have provided the scheme for the ionic fragmentation pathways. The successive work of Lopes et al.Amorim et al. (2021a) presents the PICS for 78 cations from ionization threshold to 100 eV, the PICS of the isotopologue was also measured. The absolute values of PICS for all the cations presented are in the range of 8 eV to 100 eV, only singly ionized fragments were assessed. The last work from Lopes et al.Amorim et al. (2021b) in this series of articles contains the experimental TICS of the R-carvone compared with literature data along with the theoretically predicted TICS, using the Binary encounter Bethe (BEB) Amorim et al. (2021b) model and the independent atom model (IAM) in conjunction with the screening correction additive rule (SCAR) incorporating the interference effects (I) also, which is colloquially called IAM+SCAR+I method which can be used to calculate the TICS Blanco and García (2002); Blanco and Garcıa (2003); Blanco and García (2003); Blanco et al. (2016); Chiari et al. (2013). In all the studies, Lopes et al., used Hiden analytical quadrupole mass spectrometer (QMS -EPIC 300) which can detect mass from 1 amu to 300 amu comprising a detector resolution of 1 amu which was used at residual gas analyzing (RGA) mode.
2-Butanol is a colorless organic solvent that also possesses chirality Iwahashi et al. (1999). Recently the compound has become a study of interest for many researchers as it can be a potential replacement for gasoline and can be directly used as a replacement without modifying the engines Lopes et al. (2020b) and references therein. Like other alcohols 2-butanol also belongs to the class of biofuels like methanol, ethanol, 1-proponol, and 1-butanol. Several experimental and theoretical investigations are available for these molecules Pires et al. (2018); Ghosh et al. (2018); Goswami et al. (2022) and references therein. Nirali et al.Bhavsar et al. (2023) provided a recent study of electron collision dynamics with n-butanol containing an extensive review of the available literature on n-butanol. But studies on 2-butanol are scarce. Researchers are looking to engineer microbes to produce fuels such as 1-butanol and 2-butanol and have identified an engineered strain of L. diolivorans microbe, which can be used to produce 2-butanol up to 10g/Litre from meso-2,3-butanediol through an anaerobic fermentation process Russmayer et al. (2019). Amorim et al.Amorim et al. (2022) experimentally observed the mass spectrum for 51 cations and the appearance energies for 38 cationic fragments of 2-butanol are provided in their work. For their measurements, they used the HIDEN’s energy pulse ion counting (EPIC - 300) quadruple mass spectrometer (QMS) operated on residual gas analyzing (RGA) mode Nixon et al. (2016); Pires et al. (2018); Lopes et al. (2020a). Lopes et al.provided a comprehensive review on the fragmentation of biofuels which contains all the literature till the year 2019 for the Methanol, Ethanol, Proponol, and Butanol molecules Lopes et al. (2020b). Bettega and co-workers performed low-energy electron scattering studies on isomers of butanol including 2-butanol where they provided the differential cross sections (DCS), integral cross sections (ICS), momentum transfer cross sections (MTCS) for an incident energy range of 1 eV to 50 eV Bettega et al. (2010). The preceding work of Amorim et al.provided the experimental PICS for the fragments of 2-butanol along with their experimental TICS and theoretical TICS using the BEB method and the IAM+SCAR+I method Amorim et al. (2023).
Imidazole acts as a base for several biomolecules. Imidazole and 2-nitroimidazole are very good radiosensitizers used in cancer treatment and also in the development of drugs in pharmaceutical chemistry Meissner (2019). A brief review Shalini et al. (2010); Zhang et al. (2014) of their use provides us with the importance of these targets for various applications. Rebecca et al.Meissner et al. (2019) studied the low-energy fragmentation of imidazole and 2-nitroimidazole, and has provided us with the appearance energies and mass spectrum for the cationic fragments. A recent study by Tejas et al.Jani et al. (2023) for imidazole has investigated the dissociative electron attachment (DEA), excitation, and ionization processes by electron impact.
In the present study, we have made use of the appearance energies and mass spectrum data from the literature Lopes et al. (2020a); Meissner et al. (2019); Amorim et al. (2022) for all the targets studied here. The structure of the article is as follows: In section II, we have discussed the BEB model for TICS and the computational method to calculate orbital energies required as input in the BEB model. In section II.1, various methods to calculate the PICS using the BEB model, appearance energies, and mass spectrum data are discussed. In section III, we have shown the results and their discussions. In Section IV we summarize our results and findings.
II BEB Model
The BEB model Kim and Rudd (1994) is one of the most common and widely used methods for predicting the TICS for atoms, molecules, radicals, and ions and is being employed here with some modifications to calculate the PICS. The BEB method calculates the electron impact ionization cross section for each orbital and the sum of the ionization cross section for each orbital gives the total ionization cross section [Eq. (1)] for a target under consideration. [Eq (2)] gives the BEB formula for determining the TICS.
The reduced variables and S are defined,
Here is the principal quantum number, is the Bohr radius, is Rydberg constant, is the orbital binding energy, is the orbital kinetic energies, is the orbital occupation number, is the incident electron kinetic energy and is differential oscillator strength which is set to unity in the simplified BEB model.
To calculate the orbital parameters that are required as input to the BEB model, we first optimized the target molecules using the density functional with the aug-cc-PVTZ (aVTZ) basis set, which calculates the energy minima at which the molecule is more stable. After optimization, the orbital binding and kinetic energies are calculated using the Hatree-Fock (HF) approximation that has been found to work efficiently over the years within 10 15 % of experimental uncertainty Vukstich et al. (2022). All the quantum chemistry calculations are performed using the GAUSSIAN-16 Frisch et al. (2016) quantum chemistry software. The structures of the target molecules studied are shown in Figure. 1. The main aspect of this work is to calculate the branching ratios and PICS of the possible fragments originating from the parent molecules which are described in the sections to follow.
II.1 Branching ratio (BR) and Partial ionization cross section (PICS)
When the molecular target undergoes dissociative ionization after an electron or positron impact, there will be several reaction channels in which ionization takes place. The electron impact mass spectrum (EIMS) is a characterization technique that can provide information about the various cations present, along with their mass-to-charge ratio , relative abundances, and respective appearance energies. Here, the branching ratio plays an important role, which describes the contribution of each reaction channel to the TICS. There are different methods proposed to calculate the branching ratio (BR), which in turn is used to compute the PICS. Karl IrikuraIrikura (2017) used the EIMS data from the NIST Linstrom and Mallard (2001) to calculate BR and have used that to compute the PICS of many organic molecules at a single electron impact energy. Hamilton et al.Hamilton et al. (2017) used the experimental EIMS to calculate the BR and hence the PICS, which is discussed in more detail later. Huber et al.Huber et al. (2019) proposed the way to calculate the BR using the appearance energy of the fragments by electron impact. The BR was combined with the TICS from the BEB method to obtain the PICS for various fragments which is discussed in more detail below in II.2. Hamilton et al.Hamilton et al. (2017) proposed a modification in the reduced variable , which is given in Eq. (4).
Here can either be the appearance energy (AE) or the dissociation energy for the specific fragment. The AE and ionization energy (IE) or ionization potential (IP) aren’t the same, IP is the minimum energy that is required to remove an electron from a molecule’s bound state whereas the AE is the minimum energy required to induce fragmentation of a molecule. After the modification of Eq.(4), the modified BEB formula given in Eq.(5) can be used to calculate PICS,
Using Eq.(5) the PICS of each fragment can be calculated as,
Here, is the experimental BR obtained from the electron impact mass spectrometry (EIMS). The EIMS is performed at the reference incident electron energy of 70 eV or 100 eV. The branching ratio of the fragment is defined as,
denotes the relative intensity of the particular fragment and is the total ion intensity which is the sum of all the relative intensity of the fragments that are detected. In case of unavailability of EIMS data for a molecule of interest, performing quantum chemical mass spectrometry (QCxMS) Grimme (2013) can provide information on the fragments along with their relative abundances and fragmentation pathways. The project QCxMS is very well tested Koopman and Grimme (2019) and is actively under development for more accurate prediction of the EIMS theoretically.
On the other hand Baluja et al.Goswami et al. (2021) presented a similar form of the Eq.(5), where they also call it a modified BEB method, we address their model with the superscript B making it . The key assumption in the m-BEB-B model is, that the cation that is dissociated from the neutral parent molecule can be distinguished by their , which is different from IE that corresponds to the neutral molecule. The vertical IE is the energy of the highest occupied molecular orbital (HOMO). To make the BEB model feasible across cations, a consistent alteration is made in the values of occupied orbitals binding energies. The changes are accomplished by adding the difference between IE of the neutral molecule and the of the cation to the all the values in orbital binding energy (B) of the neutral molecule (). As a result, the HOMO of the particular cation becomes its appearance energy.
The new reduced variables are,
The PICS is calculated using a similar approach used in Eq. (6)
The PICS calculated from the m-BEB-B model is then scaled using the factor , which is obtained from the experimental BR Eq.(7) and the theoretical BR which is calculated from the PICS itself. The theoretical BR for the fragment is calculated at the same incident energy as the experimental BR ,
There are other ways to calculate the theoretical BR without the need for PICS, which was introduced by Huber et al.Huber et al. (2019) which will be explained in the later part of this article. The scaling factor is obtained by
This scaling factor implemented by Baluja et al.Goswami et al. (2021) and the structural factor incorporated by Huber et al.Huber et al. (2019) and Vincent et al.Graves et al. (2022) are the same, details about structural factor are discussed in the next section.
II.2 Huber’s method
To theoretically calculate the BR of the fragments we require the information about the dissociation energies . is the branching factor, which isn’t the same as BR.
is a parameter that is used to make the slowly reach the asymptotic region and if the incident energy is less than the then is 0 as the dissociation never happens. In our case, we keep it as 3 as proposed by HuberHuber et al. (2019) The theoretical BR calculated by Huber’s method is given below and is denoted by the superscript H.
It is provided that the additional structural factor paves the way for other processes within the threshold region.
is the actual BR, and is determined from the experimental BR by taking the ratio of experimental BR to that of theoretical BR calculated using Huber’s method at the reference energy,
The way this structural factor is calculated is the same as the scaling factor which was discussed earlier. It is shown that for simple dissociation of bonds, the is taken as, and for more complex dissociation it is suggested to be Huber et al. (2019). It was considered by Vincent et al.Graves et al. (2022) for the ease of calculation of branching ratios and this could be explored further in the future.
II.3 Mass spectrum dependence (MSD) method
In all the above methods the BR is calculated at a single energy, this method was proposed to make the branching ratio dependent on the incident energy of the electron. The BR calculated using this method will be denoted with the superscript (MSD)
In this method, we would also need a BR at the reference energy, which can either be theoretical BR or experimental BR. Suppose, there is no experimental EIMS data present for the target then we can use the theoretical BR calculated from Huber’s method, referred to as the hybrid method by Vincent et al.Graves et al. (2022). This assumption is true only if we strictly interpret that the already computed branching ratios are asymptotic when the incident energy is infinity, the branching ratio reaches the known branching ratio which gives rise to the assumption that reference energy is at infinity. Also, this method does not give the same reference branching ratio at the reference incident energy Graves et al. (2022). The control parameter was taken as Janev and Reiter (2004), it was taken as 1.5 in this work as an average value. As we have discussed several methods to calculate the branching ratios, we should also remember that the sum of all the branching ratios of the fragments should sum up to give unity . However, this condition may not be satisfied if we do not have information on all the fragmentation or reaction pathways for a target molecule.
III Results and Discussions
We have theoretically calculated the PICS for all the molecular targets which we have shown in Figure. 1 using the m-BEB-B method and the MSD method. Here we did not employ Huber’s method Huber et al. (2019) and the m-BEB-A Hamilton et al. (2017) method as the magnitude of their cross sections was not as accurate as the MSD method Graves et al. (2022) or the m-BEB-B method Goswami et al. (2021) as compared to experimental results. More details can be found in the upcoming subsections.
III.1 R-Carvone
In the series of articles published by Lopes et al.Lopes et al. (2020a) and Amorim et al.Amorim et al. (2021a, b) have extensively studied the fragmentation of R-carvone by electron impact and provided the absolute PICS and TICS along with the mass spectrum and appearance energies and the Wannier exponents for 35 cations. We have calculated the theoretical PICS for these fragments using the MSD method [Eq. (17)], and mBEB-B method [Eq. (8)] based on the data presented by Lopes et al.Lopes et al. (2020a) and Amorim et al.Amorim et al. (2021a, b). The most stable cations were found to have masses of 39 amu , 54 amu , and 82 amu , as their relative abundances contribute 40% to the total ion intensity. Owing to higher contribution in the mass spectrum, their PICS also tend to have higher magnitudes than other cations which can be seen from Table.1. In the MSD method the BR is scaled for the incident energy whereas in the m-BEB-B method, the binding energies of the molecular orbitals are scaled to the appearance energy. Although they seem different from one another, but work the same when it comes to calculating the PICS. In Table.1, we have shown all the important data that we have obtained while performing our calculations. It consists of the BR obtained using Huber’s method [Eq. (14)], theoretical BR which is calculated while performing the m-BEB-B method [Eq. (11)], the experimental BR which is obtained from the EIMS data presented by Lopes et al.Lopes et al. (2020a) [Eq. (7)], the structural factor [Eq. (16)], the scaling factor [Eq. (12)] and the PICS max from MSD method and the m-BEB-B method compared with the experimental PICS max Amorim et al. (2021a). The m-BEB-B PICS presented in the table are scaled using the . We have also calculated the mean squared error (MSE) for the PICS, assuming that the experimental values are true values and the PICS we obtained as predicted values. The MSE for MSD-PICS was 0.1327 and the MSE for m-BEB-B was 0.1038 giving almost similar PICS.
In the Figure.3, It is seen that for most of the fragments, the PICS calculated using both methods shows good agreement with the experimental PICS. Some discrepancies of the present data with the experiment are seen for cations , , and where our data are found to underestimate the experimental results, the reason for the discrepancy may be due to the appearance energies and the relative abundance which are smaller in our case compared to experimental values. However, for a few cations like the PICS calculated using the MSD shows a very good agreement with the experimental data of Amorim et al., whereas for several other cations both the MSD method and the m-BEB-B method predicts the PICS within the experimental uncertainty as shown Figure.3. In Figure. 2, shows the TICS obtained using the BEB model, MSD, and m-BEB-B methods which is the sum of the PICS of all the fragments. Our results are compared with the existing experimental and theoretical data Amorim et al. (2021a, b), and the present results using the m-BEB-B method are found to be in excellent agreement with the experimental data. Our TICS using the MSD method is found to underestimate all other data because the appearance energies were not provided for all the fragments. In the series of articles Lopes et al. (2020a); Amorim et al. (2021a, b) they have presented the experimental PICS at 100 eV for 78 cationic fragments but have only provided the AE for 35 prominent cations. We believe if we could calculate the PICS for other remaining fragments and sum it then it might give a better comparison with the experimental TICS and the TICS calculated using the BEB model.
| Cations | (eV) Lopes et al. (2020a) | Lopes et al. (2020a) | Amorim et al. (2021a) | |||||||
| 14 | 15.34 | 0.033333 | 0.540558 | 0.001539 | 0.046171 | 0.002847 | 0.03263 | 0.037382 | 0.04474 | |
| 16 | O | 14.16 | 0.030769 | 0.608827 | 0.002601 | 0.084534 | 0.004271 | 0.055578 | 0.062124 | 0.05242 |
| 26 | 11.67 | 0.025358 | 0.794778 | 0.004087 | 0.161172 | 0.005142 | 0.08878 | 0.094934 | 0.08685 | |
| 28 | 12.3 | 0.026727 | 0.741232 | 0.032002 | 1.197369 | 0.043174 | 0.692441 | 0.748003 | 0.64644 | |
| 28 | 14.58 | 0.031681 | 0.583319 | 0.032002 | 1.010126 | 0.054862 | 0.682016 | 0.768848 | 0.64644 | |
| 29 | 12.04 | 0.026162 | 0.762717 | 0.014356 | 0.548735 | 0.018822 | 0.311144 | 0.334674 | 0.37457 | |
| 29 | 14.93 | 0.032442 | 0.563107 | 0.014356 | 0.442517 | 0.025494 | 0.305212 | 0.346601 | 0.37457 | |
| 31 | 11.97 | 0.026010 | 0.768646 | 0.000531 | 0.020415 | 0.000690 | 0.011508 | 0.012364 | 0.01021 | |
| 39 | 10.83 | 0.023533 | 0.874773 | 0.074061 | 3.147148 | 0.084663 | 1.617523 | 1.708448 | 2.61851 | |
| 39 | 15.61 | 0.033919 | 0.526334 | 0.074061 | 2.183447 | 0.140712 | 1.567225 | 1.806174 | 2.61851 | |
| 40 | 10.63 | 0.023098 | 0.895457 | 0.010667 | 0.461812 | 0.011913 | 0.233269 | 0.245712 | 0.2041 | |
| 41 | 12.56 | 0.027292 | 0.720556 | 0.046743 | 1.712707 | 0.06487 | 1.009673 | 1.095517 | 2.15466 | |
| 41 | 14.82 | 0.032203 | 0.569361 | 0.046743 | 1.451525 | 0.082097 | 0.994528 | 1.126763 | 2.15466 | |
| 42 | 10.32 | 0.022425 | 0.928897 | 0.005758 | 0.256772 | 0.006199 | 0.126158 | 0.132352 | 0.11159 | |
| 43 | 11.99 | 0.026053 | 0.766946 | 0.007496 | 0.287718 | 0.009774 | 0.162525 | 0.174674 | 0.15451 | |
| 45 | 11.99 | 0.026053 | 0.766946 | 0.000756 | 0.029017 | 0.000986 | 0.016396 | 0.017622 | 0.01444 | |
| 52 | 10.68 | 0.023207 | 0.890223 | 0.005387 | 0.232130 | 0.006051 | 0.117758 | 0.124124 | 0.10637 | |
| 53 | 12.52 | 0.027205 | 0.723686 | 0.024386 | 0.896380 | 0.033697 | 0.526899 | 0.571304 | 0.50587 | |
| 54 | 13.81 | 0.030008 | 0.631203 | 0.093897 | 3.129061 | 0.148758 | 2.011533 | 2.233011 | 2.43499 | |
| 55 | 12.67 | 0.027531 | 0.712042 | 0.009911 | 0.359995 | 0.013919 | 0.213932 | 0.232561 | 0.19521 | |
| 58 | 10.01 | 0.021751 | 0.964127 | 0.029481 | 1.355390 | 0.030578 | 0.647126 | 0.676302 | 0.56612 | |
| 60 | 10.62 | 0.023076 | 0.896510 | 0.000836 | 0.036227 | 0.000932 | 0.01828 | 0.019252 | 0.01595 | |
| 65 | 15.31 | 0.033267 | 0.542169 | 0.006979 | 0.209785 | 0.012872 | 0.147989 | 0.169430 | 0.1409 | |
| 67 | 12.16 | 0.026423 | 0.867692 | 0.012153 | 0.459945 | 0.016146 | 0.263205 | 0.283664 | 0.24021 | |
| 69 | 10.94 | 0.023772 | 0.863681 | 0.007894 | 0.332075 | 0.00914 | 0.172297 | 0.182264 | 0.15821 | |
| 77 | 13.12 | 0.028509 | 0.678578 | 0.010561 | 0.370448 | 0.015564 | 0.227293 | 0.249071 | 0.2091 | |
| 79 | 11.13 | 0.024185 | 0.844976 | 0.016200 | 0.669847 | 0.019172 | 0.353152 | 0.374592 | 0.32259 | |
| 80 | 9.63 | 0.020925 | 1.009962 | 0.007271 | 0.347475 | 0.007199 | 0.159963 | 0.16644 | 0.14366 | |
| 82 | 10.85 | 0.023576 | 0.872742 | 0.132679 | 5.627668 | 0.152025 | 2.897386 | 3.061113 | 2.71195 | |
| 83 | 10.9 | 0.023685 | 0.867692 | 0.007231 | 0.305301 | 0.008334 | 0.157858 | 0.166896 | 0.13467 | |
| 91 | 10.83 | 0.023533 | 0.874773 | 0.012047 | 0.511925 | 0.013772 | 0.263116 | 0.277906 | 0.23679 | |
| 92 | 9.87 | 0.021447 | 0.980661 | 0.00743 | 0.346440 | 0.007577 | 0.163231 | 0.170305 | 0.14487 | |
| 93 | 10.83 | 0.023533 | 0.874773 | 0.036261 | 1.540875 | 0.041452 | 0.791954 | 0.836472 | 0.73009 | |
| 94 | 9.72 | 0.021121 | 0.998829 | 0.007258 | 0.343642 | 0.007266 | 0.159585 | 0.166215 | 0.13833 | |
| 106 | 9.56 | 0.020773 | 1.018745 | 0.007258 | 0.349394 | 0.007124 | 0.159737 | 0.166078 | 0.3352 | |
| 107 | 9.35 | 0.020317 | 1.045762 | 0.01815 | 0.893349 | 0.017356 | 0.399984 | 0.414931 | 0.35451 | |
| 108 | 10.2 | 0.022164 | 0.942315 | 0.034881 | 1.573783 | 0.037017 | 0.764773 | 0.801114 | 0.69171 | |
| 135 | 10.27 | 0.022316 | 0.934455 | 0.003728 | 0.167056 | 0.003990 | 0.081708 | 0.085665 | 0.07175 |
III.2 2-Butanol
The PICS that we have theoretically determined for the fragments of 2-Butanol are based on the EIMS data presented by Amorim et al.Amorim et al. (2022), their experimentally provided EIMS, AEs for 51 cations. Pires et al.Pires et al. (2018) and Gosh et al.Ghosh et al. (2018) provided the EIMS, AEs, and the experimental PICS for the fragments of 1-Butanol. Goswami et al.Goswami et al. (2022) have theoretically calculated the PICS using the m-BEB-B method [Eq.(8)], which has good agreement with the existing experimental PICS presented by Pires et al.Pires et al. (2018) and the experimental TICS presented by Gosh et al.Ghosh et al. (2018). Owing to a good comparison of the m-BEB-B model with the literature data in 1-butanol, we proceed to calculate the PICS theoretically applying both the m-BEB-B model and the MSD model that requires the experimental EIMS data. In Table.2, all the relevant data required for calculating the PICS using the m-BEB-B and MSD are given similar to R-carvone. The maximum cross sections are also provided in the table using the two methods and are compared with the experimental maximum cross section. In Figure.5, we have compared our PICS data with the experimental PICS presented by Amorim et al.Amorim et al. (2023) in the entire energy range. The MSD method and m-BEB-B method for 2-butanol show excellent comparison with the experimental data of Amorim et al.Amorim et al. (2023) for most of the cationic fragments. However, for a few fragments such as , , , and , there were slight discrepancies when comparing our data with the experimental PICS which can be seen in Figure. 5. The PICS that we calculated are consistent for all cations from their AE threshold to 5000 eV, magnitudes of the PICS of the cations differ concerning their RI and AE.
Figure.4, shows the comparison of our TICS with the existing data available in the literature for 2-butanol. The present TICS calculated using the BEB method, compares very well with the experimental data and lies within the uncertainty of the experimental TICS. Also the TICS (obtained after summing all the PICS of the fragments) from the m-BEB-B and MSD methods compare well with the experiment and are within the range of uncertainty of the experimental TICS. The average experimental uncertainty is and all our present data lies within that limit.
| Cations222The forward slash denotes ”or” added for the cations with same values | (eV) Amorim et al. (2022) | Amorim et al. (2022) | Amorim et al. (2023) | |||||||
| 1 | 13.45 | 0.015970 | 0.826677 | 0.029448 | 1.843945 | 0.035623 | 0.328767 | 0.354294 | 0.0929 | |
| 2 | 15.52 | 0.010394 | 0.662681 | 0.033981 | 3.269164 | 0.051278 | 0.374218 | 0.417960 | 0.0997 | |
| 13 | 15.38 | 0.010681 | 0.672343 | 0.033674 | 3.152747 | 0.050085 | 0.371191 | 0.413481 | 0.0254 | |
| 14 | 14.84 | 0.011890 | 0.711384 | 0.032492 | 2.732773 | 0.045674 | 0.359453 | 0.396460 | 0.1330 | |
| 15 | 9.99 | 0.038974 | 1.251373 | 0.021873 | 0.561216 | 0.017479 | 0.249473 | 0.257970 | 0.594 | |
| 16 | 13.3 | 0.016517 | 0.840578 | 0.02912 | 1.763079 | 0.034643 | 0.325416 | 0.349881 | 0.0459 | |
| 17 | 10.91 | 0.029923 | 1.113007 | 0.023887 | 0.798291 | 0.021462 | 0.270959 | 0.282552 | 0.0655 | |
| 19 | 13.56 | 0.015585 | 0.816679 | 0.029689 | 1.905022 | 0.036354 | 0.331219 | 0.357553 | 0.772 | |
| 26 | 11.5 | 0.025549 | 1.035467 | 0.025179 | 0.985502 | 0.024317 | 0.284586 | 0.298677 | 0.188 | |
| 27 | 13.87 | 0.014563 | 0.789364 | 0.030368 | 2.085312 | 0.038472 | 0.338110 | 0.366797 | 0.964 | |
| 28 | 11.59 | 0.024959 | 1.024311 | 0.025376 | 1.016714 | 0.024774 | 0.286654 | 0.301163 | 0.279 | |
| 29 | 13.62 | 0.015380 | 0.811294 | 0.029821 | 1.939005 | 0.036757 | 0.332554 | 0.359333 | 0.834 | |
| 30 | 11.46 | 0.025818 | 1.040480 | 0.025091 | 0.971845 | 0.024115 | 0.283665 | 0.297576 | 0.0392 | |
| 31 | 12.91 | 0.018059 | 0.878236 | 0.028266 | 1.565195 | 0.032185 | 0.316662 | 0.338509 | 1.27 | |
| 32 | 12.5 | 0.019895 | 0.920350 | 0.027368 | 1.375620 | 0.029737 | 0.307408 | 0.326724 | 0.015 | |
| 38 | 13.11 | 0.017245 | 0.858645 | 0.028704 | 1.664470 | 0.033429 | 0.321159 | 0.344321 | 0.0289 | |
| 39 | 10.52 | 0.033375 | 1.168847 | 0.023033 | 0.690118 | 0.019706 | 0.261885 | 0.272052 | 0.198 | |
| 40 | 10.35 | 0.035047 | 1.194444 | 0.022661 | 0.646585 | 0.018972 | 0.257912 | 0.267511 | 0.0317 | |
| 41 | 12.34 | 0.020679 | 0.937536 | 0.027018 | 1.306544 | 0.028818 | 0.303780 | 0.322175 | 0.801 | |
| 42 | 10.66 | 0.032078 | 1.148352 | 0.02334 | 0.727609 | 0.020325 | 0.265148 | 0.275806 | 0.0805 | |
| 43 | 11.77 | 0.023831 | 1.002498 | 0.02577 | 1.081357 | 0.025706 | 0.290782 | 0.306157 | 0.566 | |
| 44 | 10.35 | 0.035047 | 1.194444 | 0.022661 | 0.646585 | 0.018972 | 0.257912 | 0.267511 | 0.546 | |
| 45 | 10.93 | 0.029759 | 1.110246 | 0.023931 | 0.804168 | 0.021555 | 0.271423 | 0.283094 | 2.53 | |
| 46 | 10.67 | 0.031988 | 1.146908 | 0.023362 | 0.730346 | 0.020369 | 0.265381 | 0.276075 | 0.0629 | |
| 51 | 12.33 | 0.020729 | 0.938625 | 0.026996 | 1.302309 | 0.028761 | 0.303552 | 0.321891 | 0.0118 | |
| 53 | 12.01 | 0.022431 | 0.974409 | 0.026296 | 1.172314 | 0.026986 | 0.296266 | 0.312863 | 0.0198 | |
| 54 | 9.75 | 0.041924 | 1.291528 | 0.021347 | 0.509186 | 0.016529 | 0.243820 | 0.251661 | 0.0192 | |
| 55 | 12.34 | 0.020679 | 0.937536 | 0.027018 | 1.306544 | 0.028818 | 0.303780 | 0.322175 | 0.102 | |
| 56 | 9.45 | 0.046045 | 1.344407 | 0.020691 | 0.449368 | 0.01539 | 0.236726 | 0.243829 | 0.0483 | |
| 57 | 11.38 | 0.026366 | 1.050608 | 0.024916 | 0.944997 | 0.023716 | 0.281823 | 0.295374 | 0.0974 | |
| 58 | 10.04 | 0.038395 | 1.243236 | 0.021982 | 0.572524 | 0.017682 | 0.250648 | 0.259290 | 0.0313 | |
| 59 | 10.99 | 0.029274 | 1.102023 | 0.024062 | 0.821959 | 0.021835 | 0.272813 | 0.284722 | 0.84 | |
| 60 | 10.66 | 0.032078 | 1.148352 | 0.02334 | 0.727609 | 0.020325 | 0.265148 | 0.275806 | 0.0298 | |
| 72 | 9.66 | 0.043107 | 1.307068 | 0.02115 | 0.490645 | 0.016182 | 0.241695 | 0.249305 | 0.014 | |
| 73 | 11.45 | 0.025886 | 1.041738 | 0.02507 | 0.968492 | 0.024065 | 0.283435 | 0.297300 | 0.0588 | |
| 74 | 9.98 | 0.039092 | 1.253010 | 0.021851 | 0.558970 | 0.017439 | 0.249238 | 0.257706 | 0.0196 | |
| 75 | 9.79 | 0.041412 | 1.284707 | 0.021435 | 0.517603 | 0.016685 | 0.244764 | 0.252710 | 0.00111 |
III.3 Imidazole (IMI) and 2-Nitroimidazole (2NI)
The PICS that we present here for IMI and 2NI are based on the mass spectrum and AEs presented by Rebecca et al.Meissner et al. (2019). In Table 3 and Table. 4, we have presented the BRs calculated using Huber’s method [Eq.(14)], the theoretical BRs [Eq.(11)], the experimental BRs calculated from the EIMS data [Eq.(7)] and the structural factor [Eq.(16)] and the scaling factor [(12)] along with the maximum cross sections calculated using the MSD and the m-BEB-B methods. The PICS for , , and cationic fragments of IMI is shown in Figure. 5(a) which were observed in the experiment of Rebecca et al.Meissner et al. (2019). is the cation with the lowest m/z having the second highest AE of 11.74 eV, and fragments have m/z value of 40 and same AE, & also share their AEs and their m/z value are same (41), fragment has the same m/z value of 41 but has a different AE and is the de-hydrogenated IMI which was also observed in the EIMS experiment. For fragments which has the same m/z values and AEs, the calculated PICS will be the same and is presented in one figure for both fragments. From the Figure. 5(a), it can be seen that for almost all the fragments the magnitude of the PICS from the MSD method is slightly smaller than the PICS calculated using the m-BEB-B method. Both methods match very well from the ionization threshold to 70 eV after which the cross sections show slight variation for all the fragments.
In Figure 5(b), we have compared all the TICS from the existing literature along with our data. Our BEB TICS compares very well with the BEB TICS presented by Tejas et al.Jani et al. (2023), albeit their TICS calculated using the csp-ic method has a higher magnitude in comparison to BEB data, especially at higher energies after the peak. The magnitudes of the sums of the PICS calculated using both the MSD and the m-BEB-B methods are lower than the TICS due to the lack of information on all the fragments and the AEs as our methods are dependent on the AEs and mass spectrum data of the fragments of the parent molecule.
The PICS of 2NI is shown in Figure. 6(a), which contains the PICS of 8 fragments. The smallest fragment is of m/z of 28 which has an AE of 11.44 eV and the biggest fragment is having the m/z of 97. The cation has two AEs, the fragment with the relatively lowest AE of 13.0 eV is marked with the superscript of a, and the other fragment with an AE of 16.29 eV is identified with the superscript of b, this is solely to distinguish. Like the PICS of IMI a similar trend is seen here, both the MSD PICS and the m-BEB-B PICS show a good agreement from the ionization threshold to 100 eV, after which the PICS of the MSD method falls short on comparison with the m-BEB-B PICS, but their magnitudes remain similar.
The TICS of 2NI is shown in Figure. 6(b) with the BEB model is compared with the TICS obtained after summing the cross sections of all the available fragments in the MSD and m-BEB-B methods. The TICS do not match very well with the BEB TICS, as we don’t have information on the AEs of all the fragments that appeared in the EIMS. All the details of the parameters required to compute PICS are given in the Table. 4 which contains the theoretical BRs, the experimental BRs, the scaling factors, and the structural factors along with the maximum CS.
| Cations | (eV) Meissner et al. (2019) | Meissner et al. (2019) | |||||||
| 28 | 11.74 | 0.125833 | 0.700872 | 0.112307 | 0.892509 | 0.160239 | 1.043251 | 1.128965 | |
| 40 | 14.96 | 0.060814 | 0.498982 | 0.246460 | 4.052667 | 0.493925 | 2.240973 | 2.592455 | |
| 40 | 14.96 | 0.060814 | 0.498982 | 0.246460 | 4.052667 | 0.493925 | 2.240973 | 2.592455 | |
| 41 | 11.68 | 0.127782 | 0.705604 | 0.025329 | 0.198219 | 0.035897 | 0.235377 | 0.254443 | |
| 41 | 11.68 | 0.127782 | 0.705604 | 0.025329 | 0.198219 | 0.035897 | 0.235377 | 0.254443 | |
| 41 | 14.06 | 0.073256 | 0.546657 | 0.025329 | 0.345758 | 0.046334 | 0.231722 | 0.262657 | |
| 67 | 8.76 | 0.302892 | 0.688466 | 0.007728 | 0.025516 | 0.054560 | 0.073110 | 0.378300 |
| Cations | (eV) Meissner et al. (2019) | ||||||||
| 28 | 11.44 | 0.124877 | 0.836333 | 0.112307 | 0.134285 | 0.134285 | 1.306890 | 1.388680 | |
| 30 | 11.11 | 0.136339 | 0.867858 | 0.070544 | 0.081285 | 0.081285 | 0.822306 | 0.868480 | |
| 40 | 13.74 | 0.072078 | 0.654551 | 0.246460 | 0.376533 | 0.376533 | 2.832763 | 3.155850 | |
| 56 | 11.34 | 0.128210 | 0.845716 | 0.032927 | 0.038934 | 0.038934 | 0.383366 | 0.406602 | |
| 67 | 13.00 | 0.085100 | 0.508922 | 0.037563 | 0.073808 | 0.073808 | 0.433492 | 0.475207 | |
| 67 | 16.29 | 0.043251 | 0.508922 | 0.037563 | 0.073808 | 0.073808 | 0.425538 | 0.504195 | |
| 83 | 11.12 | 0.135971 | 0.866878 | 0.038764 | 0.044717 | 0.044717 | 0.451837 | 0.477295 | |
| 97 | 13.92 | 0.069318 | 0.642637 | 0.016542 | 0.025741 | 0.025741 | 0.189945 | 0.212466 |
IV Conclusion
In the present work, we have theoretically computed the partial and total ionization cross sections of R-carvone, 2-butanol, imidazole, and 2-nitroimidazole using the MSD and m-BEB-B methods in combination with the BEB model. We obtained a fairly good agreement of our theoretical PICS data of various fragments with the experimental PICS for R-carvone and 2-butanol. Also, the TICS which is obtained as the sum of all the PICS of the fragments in the MSD and m-BEB-B method gives good agreement with TICS computed using the BEB model. For imidazole and 2-nitroimidazole, we could not find any experimental PICS study, and hence are computed in the present study for the first time. Hence, more investigation both theoretical and experimental is needed for these molecules. The appearance energies and relative abundances of the fragments were the key parameters for the determination of the PICS, which were all taken from the experimental papers for our calculations. All other parameters such as the experimental and theoretical BRs, scaling factors, and structural factors employed for the determination of PICS were computed and are presented in Tables. 1, 2, 3 and 4 respectively. The TICS data obtained after summing all the PICS of the fragments in the MSD and m-BEB-B method underestimates the TICS data computed using the BEB model for imidazole and 2-nitroimidazole. The discrepancy is due to the lack of complete information on the fragmentation of the parent molecule, where only the major dissociation fragments are given for imidazole and 2-nitroimidazole.
Acknowledgements.
S.S. acknowledges Vellore Institute of Technology for providing a research fellowship; D.G. acknowledges Science and Engineering Research Board (SERB), Department of Science and Technology (DST), Government of India (Grant No. SRG/2022/000394) for providing the financial support and computational facility.Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Author Contributions
S. Suriyaprasanth: Conceptualization (lead), Data curation (lead), Investigation (lead), Methodology (Supporting), Resources (Supporting), Validation (lead), Writing – original draft (lead), Writing – review & editing (equal); Rounak Agrawal: Data curation (supporting), Methodology (Supporting), Investigation (Supporting); Dhanoj Gupta: Conceptualization (equal), Data curation (equal), Investigation (equal), Methodology (lead), Resources (lead), Validation (equal), Writing – original draft (equal), Writing – review & editing (lead), Supervision (lead).